OpenStudy (anonymous):

X/5-x/2=3 solve the equation

OpenStudy (anonymous):

Step 1: Find a common denominator for those fractions, so you can subtract them. What does that look like?

jhonyy9 (jhonyy9):

2x-5x=30 -3x=30 x=30/(-3) x=-10